OpenStudy (anonymous):

the integral of (x/(sq root(16-x^2)) dx

OpenStudy (anonymous):

16-x^2 = t

OpenStudy (anonymous):

Use the above substitution, or you can even use 16-x^2=t^2

OpenStudy (anonymous):

2 ∫ √(16 - x²) dx = 0 let's find the antiderivative: ∫ √(16 - x²) dx = let x = 4 sin u → sin u = (x/4) → u = arcsin(x/4) dx = 4 cos u du thus, substituting, you get: ∫ √(16 - x²) dx = ∫ √[16 - (4 sin u)²] 4 cos u du = ∫ √(16 - 16 sin²u) 4 cos u du = factor out 16: ∫ √[16(1 - sin²u)] 4 cos u du = ∫ 4 √(1 - sin²u) 4 cos u du = replace (1 - sin²u) with cos²u: ∫ 4 √(cos²u) 4 cos u du = ∫ 4 cos u 4 cos u du = ∫ 16 cos²u du = recall the double-angle identity: cos²θ = [1 + cos(2θ)]/2 thus: ∫ 16 {[1 + cos(2u)]/2} du = ∫ 8 [1 + cos(2u)] du = split it and pull out the constants: 8 ∫ du + 8 ∫ cos(2u) du = 8u + 8 (1/2)sin(2u) + C = according to the double-angle identities, rewrite sin(2u) as 2sin u cos u: 8u + 8 (1/2)2sin u cos u + C = 8u + 8sin u cos u + C summing up. you have: x = 4 sin u → sin u = (x/4) → u = arcsin(x/4) hence cos u = √(1 - sin²u) = √[1 - (x/4)²] = √[1 - (x²/16)] = √[(16 - x²)/16)] = [√(16 - x²)]/4 thus, substituting back, you get: 8u + 8sin u cos u + C = 8arcsin(x/4) + 8(x/4) {[√(16 - x²)]/4} + C = 8arcsin(x/4) + (8/16) x√(16 - x²) + C = 8arcsin(x/4) + (1/2) x√(16 - x²) + C then evaluate the definite integral, yielding: 2 ∫ √(16 - x²) dx = [8arcsin(2/4) + (1/2)2√(16 - 2²)] - [8arcsin(0/4) + (1/2) 0√(16 - 0²)] = 0 8arcsin(1/2) + √(16 - 4) - (0 + 0) = 8(π/6) + √12 = (4/3)π + 2√3 (≈ 7.6528) :-) hope it helps... Bye!

OpenStudy (anonymous):

Where did you get sin and cos from?

OpenStudy (anonymous):

How do you do the Ctrl+C and Ctrl+V rohan?

OpenStudy (mimi_x3):

\[\int\limits\frac{x}{\sqrt{16-x^{2}}} dx\] let u = 16-x^2 => du/-2x \[\int\limits\frac{x}{\sqrt{u}} *\frac{du}{-2x} =>-\frac{1}{2} \int\limits\frac{1}{\sqrt{u}} du\] the integrate it..

OpenStudy (anonymous):

See Meme agrees with me, and Mimi_x3 is always right.

OpenStudy (mimi_x3):

lol, just wanted to elaborate..

OpenStudy (anonymous):

\[-\sqrt{16-x^2}\]

OpenStudy (anonymous):

you can check @Ishaan94 in yahoo answer some one asked this question 3 years ago

OpenStudy (anonymous):

2 me

OpenStudy (anonymous):

-1/2ln |16-x^2| +c ?